OpenStudy (anonymous):

You deposit $3500 in an account that pays 4.5% annual interest, compounded monthly. How long will it take for the balance to reach 5,000?

OpenStudy (anonymous):

solve \[5000=3500(1+\frac{.045}{12})^{23t}\] for t.

OpenStudy (anonymous):

typo

OpenStudy (anonymous):

\[5000=3500(1+\frac{.045}{12})^{12t}\]

OpenStudy (anonymous):

you need the steps?

OpenStudy (anonymous):

please

OpenStudy (anonymous):

divide by 3500 get \[\frac{10}{7}=(1.00375)^{12t}\]

OpenStudy (anonymous):

our answer will not be precise because .045/12 is a repeating decimal, but we will be close

OpenStudy (anonymous):

solve \[\frac{10}{7}=(1.00375)^{12t}\] via \[\ln(\frac{10}{7})=12t\ln(1.00375)\] \[t=\frac{\ln(\frac{10}{7})}{12\ln(1.00375)}\]

OpenStudy (anonymous):

now a calculator